PLEASE HELP ME>>>>>>>>
If f(x)=x^2-3x-8, then f(10)=

Answer by Alan3354(69443)

PLEASE HELP ME>>>>>>>>
If f(x)=x^2-3x-8, then f(10)=
--------------
Wherever there's an x, replace it with 10.
f(10) = 10^2 - 3*10 - 8
f(10) = 100 -30 -8
f(10) = 62
Was that hard?